1. If one glucose molecule enters into glycolysis and proceeds through the PDC and through the citric acid cycle, how many NADH molecules are generated in these reactions?

2. If one glucose molecule enters into glycolysis and proceeds through the PDC and through the citric acid cycle, how many FADH2 molecules (that can be used directly by the ETC) are generated in these reactions?

3. If one glucose molecule enters into glycolysis and proceeds through the PDC and through the citric acid cycle, how many CO2 molecules are released in these reactions?
Answer #1

glucose is converted to 2 molecules of pyruvate by glycolysis, during glycolysis 2 NADH are produced, ( during the reaction catalyzed by glyceraldehyde 3-phosphate dehydrogenase) then during the preparatory reaction, pyruvate is converted to acetyl CoA by pyruvate dehydrogenase, one Carbon dioxide is released and one NADH is produced. 2 pyruvate molecules are produced by glycolysis so the number of carbon dioxide produced by pyruvate dehydrogenase is 2, and number of NADH produced from a glucose molecule by pyruvate dehydrogenase is 2.

during the TCA cycle with one acetyl CoA, 3 NADH, 2 carbon dioxide and one FADH2 are produced. for a glucose molecule, this cycle operates twice so number of NADH produced by a glucose molecule is via TCA cycle=6, number of carbon dioxide produced from a glucose molecule via TCA cycle=4, number of FADH2 produced from a glucose molecule via TCA cycle = 2

number of NADH produced from the glucose= 2+2+6=10

number of carbon dioxide produced from glucose=2+4=6

number of FADH2 produced from glucose=2

1) number of NADH produced from the glucose= 2+2+6=10

2) number of FADH2 produced from glucose=2

3) number of carbon dioxide produced from glucose=2+4=6